# arXiv:2603.02381v1[math.AP] 02 Mar 2026

### v1

@@ 1. Introduction
In the very recent work [CD25], Cossetti and D’Arca generalized and unified many of the previous results [RS25, YZ25, AYZ25, SYZ25]. Their approach brings together the supersolution technique employed in [AS06, FS08, CAZ21, SYZ25] and the method introduced by D’Ambrosio [D’A05], thereby unifying these tools within a common framework. However, their proofs relied on the fundamental algebraic identities [CD25, Proposition 2.1 and 2.3], which hold for $p\geq 2$.
@@ 3. Main results
In this section, using the algebraic identity (2.2), we prove that [CD25, Theorems 1.1 and 1.2] can be extended to any $1<p<\infty$. We begin with the extension of [CD25, Theorem 1.1].
@@
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.2) for any $c\in\mathbb{C}$ and $p\geq 2$.
@@
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.3) for any $c\in\mathbb{C}$.
@@
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.5) for any $c\in\mathbb{C}$.
@@ 4. Applications
4.1. Extension of Cossetti and D’Arca results.
  1. (1)
  2. (2)
  3. (3)
@@
Moreover, the “virtual” extremizers of (4.1) are of the form $u=c\phi(d)$ with $c\in\mathbb{C}$ and $p\geq 2$.
Remark 4.2.
4.2. Special cases
Here, we discuss some interesting special cases of Theorem 3.3 that are not present in [CD25]. For example, let us consider an open bounded set $\Omega$ of $\mathbb{R}^{N}$ and the following eigenvalue problem:
in Theorem 3.3, we get the following general identity:
@@ 4.2. Special cases
Moreover, the “virtual” extremizers of (4.5) are of the form $u=cu_{1}$ for any $c\in\mathbb{C}$.
Remark 4.4.
@@
The above results are almost identical in structure to the general case for the $\mathcal{L}$ operator. However, we state the Euclidean ones for simplicity and to emphasize the novelty even for the standard Laplacian $\Delta$.

### v2

# Sharp remainder formulae for general weighted Hardy and Rellich type inequalities for $1<p<\infty$

@@ 1. Introduction
In the very recent work [CD25], Cossetti and D’Arca provided similar results in the form of a distributional solution to the degenerate elliptic $p$-Laplacian equation. Their approach brings together the supersolution technique employed in [AS06, FS08, RS25, CAZ21, SYZ25] and the method introduced by D’Ambrosio [D’A05], thereby unifying these tools within a common framework. However, their proofs relied on the fundamental algebraic identities [CD25, Proposition 2.1 and 2.3], which hold for $p\geq 2$.
@@ 3. Main results
In this section, using the algebraic identity (2.2), we show the extended [CD25, Theorems 1.1 and 1.2] to any $1<p<\infty$ and provide a general weighted $L^{p}$-Rellich type inequality. We begin with the extension of [CD25, Theorem 1.1].
@@
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.2) for all $c\in\mathbb{C}$ and $p\geq 2$.
@@
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.3) for all $c\in\mathbb{C}$.
@@
Here, the functions $u=c\phi$ are “virtual” extremizers of (3.5) for all $c\in\mathbb{C}$.
@@ 4. Applications
4.1. Classical implications.
In this section, we present some classical implications of Theorems 3.1, 3.3 and 3.5. The first of which is the ability to recover sharp remainder formulae of the classical $L^{p}$-Hardy inequality and its improved versions. For simplicity, in (3.2), let us take
$\sigma=\mathcal{I}_{N},\quad V=1,\quad\phi=|x|^{-\frac{N-p}{p}}\text{and}\quad Z=\frac{x}{|x|},$
then, we have the following improved $L^{p}$-Hardy type inequality:
@@
Let $1<p<\infty$. Then, for all complex-valued $u\in C_{0}^{\infty}(\mathbb{R}^{N}\backslash\{0\})$, we have
$\int_{\mathbb{R}^{N}}\left|\nabla u\cdot\frac{x}{|x|}\right|^{p}dx=\left(\frac{N-p}{p}\right)^{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{p}}dx\\ +\int_{\mathbb{R}^{N}}C_{p}\left(\nabla u,|x|^{-\frac{N-p}{p}}\nabla\left(\frac{u}{|x|^{-\frac{N-p}{p}}}\right)\cdot\frac{x}{|x|}\right)dx.$ (4.1)
Moreover, the functions $u=c|x|^{-\frac{N-p}{p}}$ are “virtual” extremizers of (4.1) for all $c\in\mathbb{C}$ and $p\geq 2$.
Different versions of this kind of identity (4.1) have already appeared in many different works (see, e.g., [LLT21, KY24, YZ25]). The same choice of functions $V$ and $\phi$, in (3.3), gives the following $L^{p}$-Hardy identity for complex-valued functions [CKL+24, Lemma 3.3]:
@@
Let $1<p<\infty$. Then, for all complex-valued $u\in C_{0}^{\infty}(\mathbb{R}^{N}\backslash\{0\})$, we have
$\int_{\mathbb{R}^{N}}|\nabla u|^{p}dx=\left(\frac{N-p}{p}\right)^{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{p}}dx+\int_{\mathbb{R}^{N}}C_{p}\left(\nabla u,|x|^{-\frac{N-p}{p}}\nabla\left(\frac{u}{|x|^{-\frac{N-p}{p}}}\right)\right)dx.$ (4.2)
Moreover, the functions $u=c|x|^{-\frac{N-p}{p}}$ are “virtual” extremizers of (4.2) for all $c\in\mathbb{C}$.
@@
The above results are almost identical in structure to the general case for the $\mathcal{L}$ operator. However, we state the Euclidean ones for simplicity and to emphasize the novelty even for the standard Laplacian $\Delta$ in the case of the Rellich inequality.
4.2. Extension of Cossetti and D’Arca results and more.
  1. (1)
  2. (2)
  3. (3)
Corollary 4.8.
Moreover, the functions $u=c\phi(d)$ are “virtual” extremizers of (4.5) for all $c\in\mathbb{C}$ and $p\geq 2$.
Remark 4.9.
There is also one more interesting application of Theorem 3.3, which is not present in [CD25]. To be more precise, let us consider an open bounded set $\Omega$ of $\mathbb{R}^{N}$ and the following eigenvalue problem:
in Theorem 3.3, we get the following general identity:
Corollary 4.10.
Moreover, the functions $u=cu_{1}$ are “virtual” extremizers of (4.9) for all $c\in\mathbb{C}$.
Remark 4.11.